# Complex Saddles of Charged-AdS Gravitational partition function

### v1

@@ II.1.1 Complex saddles and their relevance for $\beta>\beta_{\rm max}$
After defining the starting integration contour $\mathbb{D}\in[0,\infty)$ and determining all the saddles ($z_{\sigma}$) of eq. (11), we proceed to find the relevant saddles which contribute to the integral utilizing the Picard-Lefschetz method. According to this method, we analytically continue the real $r_{+}$ to a complex variable $r_{+}=r_{+}^{x}+i\,r_{+}^{y}$. When $r_{+}$ is complex, the exponent $\mathcal{I}(r_{+})$ becomes a holomorphic function of $r_{+}$ with real ($\mathfrak{h}$) and imaginary ($\mathcal{H}$) part: $\mathcal{I}(r_{+})=\mathfrak{h}(r_{+})+i\mathcal{H}(r_{+})$. The direction of steepest descent/ascent ($\mathcal{J}_{\sigma}/\mathcal{K}_{\sigma}$) in the complex $r_{+}$ plane are given by the equation
@@
where $\lambda$ is the real parameter along the curve. - sign corresponds to the steepest descent and ++ sign corresponds to the steepest ascent. Along these curves $\mathcal{H}$ remains constant. i.e., $d\mathcal{H}/d\lambda=0$. At the saddles, one satisfies $\partial r_{+}^{x}/\partial\lambda=\partial r_{+}^{y}/\partial\lambda=0$. Note that as $\mathcal{I}(r_{+})$ is purely real when $r_{+}$ is real, the Morse function satisfies $\mathfrak{h}(r_{+}^{x},-r_{+}^{y})=\mathfrak{h}(r_{+}^{x},r_{+}^{y})$ and hence the flowlines are invariant under $(r_{+}^{x},r_{+}^{y})\rightarrow(r_{+}^{x},r_{+}^{y})$. In the complex $r_{+}$-plane, the flowlines are symmetric around the real axis ($r_{+}^{y}=0$), see appendix B for more details. The real integration contour ($\mathbb{D}$) can be deformed to a new contour ($\mathcal{C}$), which is the summation of steepest contours ($\mathcal{J}_{\sigma}$)
@@
which can never vanish for arbitrary dimensions. It is easy to convince oneself that the same will hold for an arbitrary $\beta>\beta_{\rm max}$, by observing that the integral along the real axis is absolutely convergent, and hence the steepest ascent flow-line of any saddle not lying on the real line can never intersect the real line. Hence, $n_{\sigma}$ for complex saddles is always zero making such complex saddles irrelevant in the computation of partition function. Hence, the partition function receives contributions only from the thermal-AdS for $\beta>\beta_{\rm max}$ thereby giving
@@
This result reveals two important facts: one recovers the AdS-CFT conjecture and the standard thermodynamic perspective, according to which thermal ads should only contribute to the partition function. Secondly, since the complex saddles are naked singularities and don’t contribute to the partition function for any value of $\beta>\beta_{\rm max}$, despite being dominant saddles for certain regimes of $\beta$, this shows that the results are compatible with the cosmic censorship hypothesis, which states that naked singularities are unphysical. Indeed, Picard-Lefschetz cleanly discards these complex saddles even though they are solutions of Einstein’s equation and are even dominant saddles for certain regimes of $\beta$.
@@ II.2 $Z(\beta)$ for $\beta\leq\beta_{\rm max}$
In the high temperature regime (when $\beta\leq\beta_{\rm max}$), all the saddles are purely real as can be seen from eq. (15), consequently $\theta$ is also real. This gives
@@
This implies that the action $\mathcal{I}(r_{+})$ at various saddles is also real as can be noticed from eq. (18). As range of $\beta$ is $0\leq\beta\leq\beta_{\rm max}$, this implies that $0\leq\theta<\infty$ ($\theta\to\infty$ correspond to $\beta\rightarrow 0$. From eq. (18) it can be immediately inferred that we have $\mathcal{I}(r_{+}^{-})<0$. The action at saddle $r_{+}^{+}$ is always greater than the action at saddle $r_{+}^{-}$: $\mathcal{I}(r_{+}^{-})<\mathcal{I}(r_{+}^{+})$. $\mathcal{I}(r_{+}^{+})$ changes sign depending on $\beta$ and dimensions. If $\tanh\theta<1/(d-1)$, then $\mathcal{I}(r_{+}^{+})<0$; and for $\tanh\theta>1/(d-1)$ one has $\mathcal{I}(r_{+}^{+})>0$. In the language of $\beta$ this means that for $\beta<2l\pi/(d-1)$, we have $\mathcal{I}(r_{+}^{+})<0$, while for $\beta>2l\pi/(d-1)$ the action $\mathcal{I}(r_{+}^{+})>0$. This is the Hawking-Page transition Hawking:1982dh . Summarizing it, we have
$\mathcal{I}(r_{+}^{-})<0\,\hskip 14.22636pt\forall\,\,0\leq\beta\leq\beta_{\rm max}$
$\mathcal{I}(r_{+}^{+})\leq 0\hskip 14.22636pt\forall\,\,0\leq\beta\leq\frac{2l\pi}{d-1}$
@@
For these saddles, the action $\mathcal{I}(r_{+})$ being always real implies that their imaginary part trivially vanishes. This means that the steepest ascent emanating from them always intersects our original integration contour $\mathbb{D}$, which lies in the range $0\leq r_{+}<\infty$. If we now look at the imaginary part of ${\rm Im}(\mathcal{I}(r_{+}^{+}))$, which we define as $\mathcal{H}$, we notice that $\mathcal{H}=0$ for all the saddles, the saddles lie on a Stokes ray. In such situations where the steepest ascent thimble of a saddle intersects the steepest descent thimble of another saddle, the Picard-Lefschetz method breaks down. To find out $n_{\sigma}$ unambiguously and hence the relevance of saddles, one needs to resolve such degeneracy. As the Stokes ray lies on the real axis ($r_{+}^{y}=0$), which is the axis of symmetry of the flowlines, to resolve the Stokes ray, one needs to break the symmetry; see Appendix B for more details. There are various ways one can break the symmetry, for example, by complexifying $\beta$, as in Maldacena:2019cbz . It can also be achieved by complexifying Newton’s constant ($G=|G|e^{i\epsilon},|\epsilon|\ll 1$)Ailiga:2025fny ; Mahajan:2025bzo . Choosing either sign of $\epsilon$, we find that all three saddles are relevant by the Picard-Lefschetz, see fig. (3). We also note that such complexification of parameters often modifies the asymptotic convergence, and it could be a threat to the convergence of the integral in general minisuperspace (as seen in Kanazawa:2014qma ; Ailiga:2025fny ). However, in the present case, as long as $|\epsilon|\ll 1$, the integral in eq. (11) remains convergent and well-defined.
@@
where $i$ refers to the saddles participating in the Stokes jump. $\sigma_{ij}$ and $\eta_{ij}$ are identically zero whenever there is no Stokes jump. Also, $\sigma_{ii}=\eta_{ii}=0$, and they are elements of nilpotent matrices. Whenever Stokes jump happens, smoothness across the jump requires
@@
To be concrete, let’s consider the example of the AdS Schwarzschild black hole. In this case, we observed Stokes ray among thermal a.d.s, small black hole (s.b.h) and large black hole (l.b.h) saddles when $\beta<\beta_{\rm max}$. To proceed, let us fix the orientation of the thimbles: we take the orientation to be positive (+1+1) from left to right for $\mathcal{J}(\rm a.d.s)$ and $\mathcal{J}(\rm l.b.h)$. And from down to up for $\mathcal{J}(\rm s.b.h)$. For the reverse orientation, we assign (1)(-1). With this convention, the change in orientation of the thimble as we tune $\epsilon$ from positive to negative (see fig. 3), can be encoded in the matrix form as
@@
The matrix in the above equation encodes all the elementary jumps of each thimble. The jump in eq. (29) is also expected by realizing that Morse function $\mathfrak{h}(\rm s.b.h)$ is always subdominant that $\mathfrak{h}(\rm a.d.s)$ and $\mathfrak{h}(\rm l.b.h)$. However, the dominance between $\mathfrak{h}(\rm a.d.s)$ and $\mathfrak{h}(\rm l.b.h)$ changes at Hawking-Page transition. Since the integral over $\mathcal{J}(\rm s.b.h)$ is exponentially smaller than both $\mathcal{J}(\rm a.d.s)$, and $\mathcal{J}(\rm l.b.h)$, the thimble $\mathcal{J}(\rm s.b.h)$ can’t jump multiple of $\mathcal{J}(\rm a.d.s)$, and $\mathcal{J}(\rm l.b.h)$. Also, the jump matrix remains unchanged across the transition. Comparing the transformation in eq. (29) with eq. (27), we get the jump matrix for the thimbles ($\mathcal{J}_{\sigma}$)
@@
To ensure the smoothness across the jump, i.e., to ensure $n_{\sigma}(\rm a.d.s)\mathcal{J}(a.d.s)+n_{\sigma}(\rm s.b.h)$ $\mathcal{J}(\rm s.b.h)+n_{\sigma}(\rm l.b.h)\mathcal{J}(l.b.h)$ to be continuous across the jump, one requires the condition eq. (28) to hold. From it, one can compute the jump matrix for the intersection numbers ($n_{\sigma}$), which reads
@@
Remembering that the integration contour was initially defined from 0 to $\infty$, one should put $\frac{1}{2}\mathcal{J}(\rm a.d.s)$ in eq. (33) to account only half of the thimble, Mahajan:2025bzo ; Singhi:2025rfy . For $d>4$, we encounter a degeneracy of order (d3)(d-3) for the thermal a.d.s saddle. The presence of degeneracy for a.d.s saddle wouldn’t spoil the above analysis, but demands additional care. The appearance of degeneracy implies the existence of multiple steepest-descent thimbles ($\mathcal{J}(\rm a.d.s)$) emanating from the saddle (precisely (d2)(d-2) directions, as explained below in sec. II.2.1). Only the part of the thimble lying on the positive-real axis participates in the Stokes jump, and the other thimbles show no qualitative change (see fig. 3). However, eq. (33) is still applicable, as the contribution from the extra thimbles automatically drops out while moving along the flowline. Hence, one can use the same argument and take $\frac{1}{2}\mathcal{J}(\rm a.d.s)$. Since $n_{\sigma}$ is computed using $\rm Int(\mathbb{D},\mathcal{K}_{\sigma})$ and $n_{\sigma}$ remains unaffected for a.d.s and l.b.h saddles in the homology jump, $\mathcal{K}_{\sigma}$’s corresponding to these saddles don’t change. Now, since the integration over each of the thimbles $\mathcal{J}$’s are convergent, in the saddle point approximation, one can compute the partition function eq. (11) in the high temperature limit,
@@
From eq. (34), we note that the partition function is different for $\epsilon>0$ and $\epsilon<0$ due to the sign difference in the second term and also not real. However, in the limit of vanishing $\epsilon$ (Real $G$), one gets the real answer from either side. These can be done systematically by taking the “homology average” of these two contours. By taking such averaging, one drops the contribution from the subdominant saddle Mahajan:2025bzo , and one gets
$Z(\beta<\beta_{\rm max})=\frac{1}{2}\int_{\mathcal{J}(\rm a.d.s)}dr_{+}\exp(\mathcal{I}(r_{+}))+\int_{\mathcal{J}(\rm l.b.h)}dr_{+}\exp(\mathcal{I}(r_{+})).$ (35)
@@
When $\beta=\beta_{\rm max}$, one encounters Stokes ray between thermal ads ($r_{+}=0$) and the degenerate saddle ($r_{+}=r_{+}^{+}=r_{+}^{-}$). In this case, one also observes a similar homology jump in the thimbles, see fig (4). Having obtained the intersection numbers($n_{\sigma}$), the partition function can be computed in a straightforward way by going to the cubic order (see, Feldbrugge:2017kzv for a related scenario).
As one crosses $\beta=\beta_{\rm max}$, one also observes the homological jump in the Lefschetz thimbles structure as the nature of saddles changes. Such situations are often referred to as singularities, and the Stokes ray always passes through them. To see how the thimble structure changes as we move through the singularity, we define a new variable $\mathcal{\varepsilon}=1-\frac{\beta}{\beta_{\rm max}}$, following Witten:2010cx (sec. 3.4). The critical points becomes near $\beta=\beta_{\rm max}$
$r_{+}=0,\cdots,0\,;\quad r_{+}^{\pm}=\frac{2\pi l^{2}}{\beta_{\rm max}d}\pm\frac{2\sqrt{2}\pi l^{2}\sqrt{\varepsilon}}{\beta_{\rm max}d}$ (36)
The change in the thimble as one switches from $\varepsilon<0$ to $\varepsilon>0$, the homology jump can be expressed as
@@
II.2.1 Direction of steepest descent near saddles

### v2

@@ II.1.1 Complex saddles and their relevance for $\beta>\beta_{\rm max}$
After defining the starting integration contour $\mathbb{D}\in[0,\infty)$ and determining all the saddles of eq. (11), we proceed to find the relevant saddles which contribute to the integral utilizing the Picard-Lefschetz method. According to this method, we analytically continue the real $r_{+}$ to a complex variable $r_{+}=r_{+}^{x}+i\,r_{+}^{y}$. When $r_{+}$ is complex, the exponent $\mathcal{I}(r_{+})$ becomes a holomorphic function of $r_{+}$ with real ($\mathfrak{h}$) and imaginary ($\mathcal{H}$) part: $\mathcal{I}(r_{+})=\mathfrak{h}(r_{+})+i\mathcal{H}(r_{+})$. The direction of steepest descent/ascent ($\mathcal{J}_{\sigma}/\mathcal{K}_{\sigma}$) in the complex $r_{+}$ plane are given by the equation
@@
where $\lambda$ is the real parameter along the curve. In the above equation, - sign corresponds to the steepest descent and ++ sign corresponds to the steepest ascent. Along these curves $\mathcal{H}$ remains constant. i.e., $d\mathcal{H}/d\lambda=0$. At the saddles, one satisfies $\partial r_{+}^{x}/\partial\lambda=\partial r_{+}^{y}/\partial\lambda=0$. Note that as $\mathcal{I}(r_{+})$ is purely real when $r_{+}$ is real, the Morse function satisfies $\mathfrak{h}(r_{+}^{x},-r_{+}^{y})=\mathfrak{h}(r_{+}^{x},r_{+}^{y})$ and hence the flowlines are invariant under $(r_{+}^{x},r_{+}^{y})\rightarrow(r_{+}^{x},-r_{+}^{y})$. In the complex $r_{+}$-plane, the flowlines are symmetric around the real axis ($r_{+}^{y}=0$), see appendix B for more details. The real integration contour ($\mathbb{D}$) can be deformed to a new contour ($\mathcal{C}$), which is the summation of steepest contours ($\mathcal{J}_{\sigma}$)
@@
which can never vanish for arbitrary dimensions. It is easy to convince oneself that the same will hold for an arbitrary $\beta>\beta_{\rm max}$, by observing that the integral along the real axis is absolutely convergent, and hence the steepest ascent flow-line of any saddle not lying on the real line can never intersect the real line. This implies, $n_{\sigma}$ for complex saddles is always zero making such complex saddles irrelevant in the computation of partition function (see fig. 2). Hence, the partition function receives contributions only from the thermal-AdS for $\beta>\beta_{\rm max}$ thereby giving
@@
This result reveals two important facts: Firstly, one recovers the AdS-CFT conjecture and the standard thermodynamic perspective, according to which thermal AdS should only contribute to the partition function. Secondly, since the complex saddles are naked singularities and don’t contribute to the partition function for any value of $\beta>\beta_{\rm max}$, despite being dominant saddles for certain regimes of $\beta$, this shows that the results are compatible with the cosmic censorship hypothesis Penrose:1964wq , which states that naked singularities are unphysical. Indeed, Picard-Lefschetz cleanly discards these complex saddles even though they are solutions of Einstein’s equation and are even dominant saddles for certain regimes of $\beta$.
Figure 2: Steepest descent/ascent lines are plotted in black/red for $AdS_{6}$, for $\beta=5\beta_{\rm max}$, $l=1$. Thermal-AdS is at $r_{+}=0$ and $r_{+}^{\pm}$ are naked singularities. The purple arrow is the direction of integration cycle. As the steepest ascent curves from these complex saddles don’t intersect the real line, hence they don’t contribute to the partition function.
@@ II.2 $Z(\beta)$ for $\beta\leq\beta_{\rm max}$
In the high temperature regime (when $\beta\leq\beta_{\rm max}$), all the saddles are purely real which can be seen from eq. (15), Consequently, $\theta$ is also real. This gives
@@
This implies that the action $\mathcal{I}(r_{+})$ at various saddles is also real as can be noticed from eq. (18). As range of $\beta$ is $0\leq\beta\leq\beta_{\rm max}$, this implies that $0\leq\theta<\infty$ ($\theta\to\infty$ correspond to $\beta\rightarrow 0$). From eq. (18) it can be immediately inferred that we have $\mathcal{I}(r_{+}^{-})<0$. The action at saddle $r_{+}^{+}$ is always greater than the action at saddle $r_{+}^{-}$: $\mathcal{I}(r_{+}^{-})<\mathcal{I}(r_{+}^{+})$. $\mathcal{I}(r_{+}^{+})$ changes sign depending on $\beta$ and dimensions. If $\tanh\theta<1/(d-1)$, then $\mathcal{I}(r_{+}^{+})<0$; and for $\tanh\theta>1/(d-1)$ one has $\mathcal{I}(r_{+}^{+})>0$. In the language of $\beta$ this means that for $\beta<2l\pi/(d-1)$, we have $\mathcal{I}(r_{+}^{+})<0$, while for $\beta>2l\pi/(d-1)$ the action $\mathcal{I}(r_{+}^{+})>0$. This is the Hawking-Page transition Hawking:1982dh . Summarizing it, we have
$\mathcal{I}(r_{+}^{-})<0\,\hskip 14.22636pt\forall\,\,0\leq\beta\leq\beta_{\rm max}\,$,
$\mathcal{I}(r_{+}^{+})\leq 0\hskip 14.22636pt\forall\,\,0\leq\beta\leq\frac{2l\pi}{d-1}\,$,
@@
In the next subsequent section, we analyse the partition function stated in eq. (11) using Picard-Lefschetz methods for the case of $\beta<\beta_{\rm max}$.
II.2.1 Real saddles and their relevance for $\beta<\beta_{\rm max}$
For the real saddles, the action $\mathcal{I}(r_{+})$ being always real implies that their imaginary part trivially vanishes. This means that the steepest ascent emanating from them always intersects our original integration contour $\mathbb{D}$, which lies in the range $0\leq r_{+}<\infty$. If we now look at the imaginary part of ${\rm Im}(\mathcal{I}(r_{+}^{+}))$, which we define as $\mathcal{H}$, we notice that $\mathcal{H}=0$ for all the saddles, the saddles lie on a Stokes ray. In such situations where the steepest ascent thimble of a saddle intersects the steepest descent thimble of another saddle the Picard-Lefschetz method breaks down and the associated flowline is known as Stokes ray. To find out $n_{\sigma}$ unambiguously and hence the relevance of saddles, one needs to resolve such degeneracy. As the Stokes ray lies on the real axis ($r_{+}^{y}=0$), which is the axis of symmetry of the flowlines, to resolve the Stokes ray, one needs to break the symmetry; see appendix B for more details. There are various ways one can break the symmetry, for example, by complexifying $\beta$, as in Maldacena:2019cbz . It can also be achieved by complexifying Newton’s constant ($G=|G|e^{i\epsilon},|\epsilon|\ll 1$)Ailiga:2025fny ; Mahajan:2025bzo ; Ivo:2025yek . Choosing either sign of $\epsilon$, we find that all three saddles are relevant by the Picard-Lefschetz, see fig. (3). We also note that such complexification of parameters often modifies the asymptotic convergence, and it could be a threat to the convergence of the integral in general minisuperspace (as seen in Kanazawa:2014qma ; Ailiga:2025fny ). However, in the present case, as long as $|\epsilon|\ll 1$, the integral in eq. (11) remains convergent and well-defined.
@@ II.2.1 Real saddles and their relevance for $\beta<\beta_{\rm max}$
where $i$ refers to the saddles participating in the jump. $\sigma_{ij}$ and $\eta_{ij}$ are identically zero whenever there is no Stokes jump. Also, $\sigma_{ii}=\eta_{ii}=0$, and they are elements of nilpotent matrices. Whenever Stokes jump happens, smoothness across the jump requires
@@
To be concrete, let’s consider the example of the AdS Schwarzschild black hole. In this case, we observed Stokes ray among thermal a.d.s, small black hole (s.b.h) and large black hole (l.b.h) saddles when $\beta<\beta_{\rm max}$. To proceed, let us fix the orientation of the thimbles: we take the orientation to be positive (+1+1) from left to right for $\mathcal{J}(\rm a.d.s)$ and $\mathcal{J}(\rm l.b.h)$, while from down to up for $\mathcal{J}(\rm s.b.h)$. For the reverse orientation, we assign (1)(-1). With this convention, the change in orientation of the thimble as we tune $\epsilon$ from positive to negative (see fig. 3), can be encoded in the matrix form as
@@
The matrix in the above equation encodes all the elementary jumps of each thimble. The jump in eq. (29) is also expected by realizing that Morse function $\mathfrak{h}(\rm s.b.h)$ is always subdominant than $\mathfrak{h}(\rm a.d.s)$ and $\mathfrak{h}(\rm l.b.h)$. However, the dominance between $\mathfrak{h}(\rm a.d.s)$ and $\mathfrak{h}(\rm l.b.h)$ changes at Hawking-Page transition. Since the integral over $\mathcal{J}(\rm s.b.h)$ is exponentially smaller than both $\mathcal{J}(\rm a.d.s)$, and $\mathcal{J}(\rm l.b.h)$, the thimble $\mathcal{J}(\rm s.b.h)$ can’t jump multiple of $\mathcal{J}(\rm a.d.s)$, and $\mathcal{J}(\rm l.b.h)$. Also, the jump matrix remains unchanged across the transition. Comparing the transformation in eq. (29) with eq. (27), we get the jump matrix for the thimbles ($\mathcal{J}_{\sigma}$)
@@
To ensure the smoothness across the jump, i.e., to ensure $n_{\sigma}(\rm a.d.s)\mathcal{J}(a.d.s)+n_{\sigma}(\rm s.b.h)$ $\mathcal{J}(\rm s.b.h)+n_{\sigma}(\rm l.b.h)\mathcal{J}(l.b.h)$ remains continuous across the jump, one requires the condition eq. (28) to hold. From it, one can compute the jump matrix for the intersection numbers ($n_{\sigma}$), which reads
@@
Remembering that the integration contour was initially defined from 0 to $\infty$, one should put $\frac{1}{2}\mathcal{J}(\rm a.d.s)$ in eq. (33) to account only half of the thimble, Mahajan:2025bzo ; Singhi:2025rfy . For $d>4$, we encounter a degeneracy of order (d3)(d-3) for the thermal a.d.s saddle. The presence of degeneracy for a.d.s saddle wouldn’t spoil the above analysis, but demands additional care. The appearance of degeneracy implies the existence of multiple steepest-descent thimbles ($\mathcal{J}(\rm a.d.s)$) emanating from the saddle (precisely (d2)(d-2) directions, as explained below in sec. II.2.2). Only the part of the thimble lying on the positive-real axis participates in the Stokes jump, and the other thimbles show no qualitative change (see fig. 3). However, eq. (33) is still applicable, as the contribution from the extra thimbles automatically drops out while moving along the flowline. Hence, one can use the same argument and take $\frac{1}{2}\mathcal{J}(\rm a.d.s)$. Since $n_{\sigma}$ is computed using $\rm Int(\mathbb{D},\mathcal{K}_{\sigma})$ and $n_{\sigma}$ remains unaffected for a.d.s and l.b.h saddles in the homology jump, $\mathcal{K}_{\sigma}$’s corresponding to these saddles don’t change. Now, since the integration over each of the thimbles $\mathcal{J}$’s are convergent, in the saddle point approximation, one can compute the partition function eq. (11) in the high temperature limit,
@@
From eq. (34), we note that the partition function is different for $\epsilon>0$ and $\epsilon<0$ due to the sign difference in the second term and also not real. However, in the limit of vanishing $\epsilon$ (Real $G$), one gets the real answer from either side. These can be done systematically by taking the “homology average” of these two contours. By taking such averaging, one drops the contribution from the subdominant saddle Mahajan:2025bzo , and gets
$Z(\beta<\beta_{\rm max})=\frac{1}{2}\int_{\mathcal{J}(\rm a.d.s)}dr_{+}\exp(\mathcal{I}(r_{+}))+\int_{\mathcal{J}(\rm l.b.h)}dr_{+}\exp(\mathcal{I}(r_{+}))\,$, (35)
@@
where only thermal a.d.s and large black hole contribute to the partition function in the high temperature limit. When $\beta=\beta_{\rm max}$, one encounters Stokes ray between thermal ads ($r_{+}=0$) and the degenerate saddle ($r_{+}=r_{+}^{+}=r_{+}^{-}$). In this case, one also observes a similar homology jump in the thimbles, see fig (4). Having obtained the intersection numbers($n_{\sigma}$), the partition function can be computed in a straightforward way by going to the cubic order Feldbrugge:2017kzv (see, sec. II.2.2 for detailed discussion).
As one crosses $\beta=\beta_{\rm max}$, one also observes the homological jump in the Lefschetz thimbles structure as the nature of saddles changes. Such situations are often referred to as singularities, and the Stokes ray always passes through them. To see how the thimble structure changes as we move through the singularity, we define a new variable $\mathcal{\varepsilon}=1-\frac{\beta}{\beta_{\rm max}}$, following Witten:2010cx (sec. 3.4). The critical points become near $\beta=\beta_{\rm max}$
$r_{+}=0,\cdots,0\,;\quad r_{+}^{\pm}=\frac{2\pi l^{2}}{\beta_{\rm max}d}\pm\frac{2\sqrt{2}\pi l^{2}\sqrt{\varepsilon}}{\beta_{\rm max}d}\,.$ (36)
The orientation of the thimble changes as one switches from $\varepsilon<0$ to $\varepsilon>0$ . The resultant homology jump can be expressed as
@@
II.2.2 Direction of steepest descent near saddles